OpenStudy (anonymous):

Annmarie can plow a field in 240 minutes. Gladys can plow a field 80 minutes faster. If they work together, how many minutes does it take them to plow the field?

OpenStudy (anonymous):

@amoodarya

OpenStudy (amoodarya):

\[\frac{ A }{ 240 }=\frac{ G }{240-80 }\\A=\frac{3}{4}G\]

OpenStudy (anonymous):

so it would be 160?

OpenStudy (amoodarya):

you have to find A+G

OpenStudy (anonymous):

how?

OpenStudy (amoodarya):

oh . sorry i had a mistake

OpenStudy (anonymous):

SO WHAT WE DO?

OpenStudy (amoodarya):

\[\frac{ a }{ 240 }=\frac{ g }{160 }\\\frac{ a }{ 24 }=\frac{ g }{16 }\\\frac{ a }{ 3 }=\frac{ g }{2 }\]

OpenStudy (anonymous):

then?

OpenStudy (phi):

I would use the idea of rate * time = amount of work rate is how much work you do per minute for example A does 1 field/240 minutes B goes 80 minutes faster so 1 field/160 minutes you can add the rates (1/240 + 1/160)* t = 1 field add the two fractions, and then find t

OpenStudy (amoodarya):

\[\frac{ a }{ 3 }=\frac{ g }{ 2 }=\frac{ a+g }{x }\] find x

OpenStudy (amoodarya):

"phi" said it very simpler ,and clear

OpenStudy (anonymous):

96?

OpenStudy (phi):

yes, 96 minutes

OpenStudy (anonymous):

thank u guys